A regular furnace costs $900. A high-efficiency furnace costs $1500. The average monthly bill with the regular furnace is expected to be $80. The average monthly bill with the high-efficiency furnace is expected to be $50. How long until the high-efficiency furnace pays for itself? Assume an interest rate of 1% per month
Answer #1

For deciding the number of months till high efficiency furnace pays for itself.

Cash flow associated with regular furnace = cash flow associated with high efficiency furnace.

Remember that present discounted value (PV) of any future cash flow (FV) at interest rate r received at the end of nth month is given by -

PV= (1+r)

50 50 80 80 1500+t1.01)m 80 900+ 1.011.012 t (1,01ア (1.01)n

80- 50 80-50 80-50 1.01 T (10112 +, + ( 1 , 01)-= 600

30 30 0 (1.01)2 1.010-600

1.01 (1.01)21.01600/30 20

Sum of a finite G.P. with n terms - first term "a" and common ratio "r" is given by (when r < 1)

a1- 0.99(1- 0.99) S= = 99(1-0.99) 0.01 where r = 1/1.01 = 0.99

Hence, equating LHS with RHS, we get:

99(1-0.99) = 20 1-0.99-20/99 0.202→ 0.99-1-0.202

0.99-0.798nln(0.99) In(0.798)

n = ln(0.798)/ln(0.99) = 22.45

Hence by the end of 22 months, high efficiency furnace would have paid it for itself.